OpenStudy (anonymous):

integrate ((square root of x) +2)/5(square root of x)

OpenStudy (anonymous):

\[\frac{ \sqrt{x}+2 }{ 5\sqrt{x} }\] Like this?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

@shiraz14

OpenStudy (anonymous):

First thing you want to do is factor out the 1/5

OpenStudy (anonymous):

\[\frac{ 1 }{ 5 }\int\limits \frac{ \sqrt{x}+2 }{ \sqrt{x} }dx\]

OpenStudy (anonymous):

\[u = \sqrt{x} \] \[du = \frac{ 1 }{ 2\sqrt{x} }dx\] \[\frac{ 2 }{ 5 }\int\limits (u+2)du\]

OpenStudy (anonymous):

Can you finish it off?

OpenStudy (anonymous):

You should get \[\frac{ 1 }{ 5 }(x+4\sqrt{x})+C\]

OpenStudy (anonymous):

@mpriole Did you understand what I did, came back to make sure you understood the problem :)

OpenStudy (anonymous):

what happened when you multiplied by ?i see that the 2 went out front but what happened to the square root of x?